# From Forecasting to Planning: Policy World Model for Collaborative State-Action Prediction

### v1

@@ 1 Introduction
Driving world models have recently garnered growing research interest, due to their capacity to simulate future environmental states, enabling autonomous systems to anticipate complex traffic dynamics and enhance decision-making safety [1, 2]. This trend is particularly evident in video-based paradigms, driven by the accessibility of large-scale video datasets and the breakthroughs of the video generation technique [3, 4, 5]. Despite remarkable progress having been achieved, existing world model based autonomous driving methods  [6, 7, 8, 9, 10] mostly operate in a decoupled manner (Figure 1 (a)), i.e., the world models aim to predict the next state and the associated reward but cannot directly perform trajectory planning. A separate policy model is still required for perceiving and planning. As a result, the full potential of the world model in autonomous driving is largely restricted [11].
Several concurrent works have made the initial attempt towards integrating both world modeling and planning in a unified autoregressive model [12, 11], where interleaved image and action token sequences are generated through the next-token prediction manner. Although the two tasks are jointly learned in these works, they are still independently conducted (Figure 1 (b)), where world modeling prioritizes high-fidelity next frame prediction with photo-realistic details conditioned on input actions, and trajectory planning is performed through an end-to-end mapping from the visual observation to output actions without explicitly leveraging the learned world model. As such, the world modeling and trajectory planning are only unified in terms of model architecture and prediction manner, while their synergistic mechanism remains under-explored. It is still unknown whether and how this unification can further benefit autonomous driving.
@@
Figure 1: Comparison of video world models for autonomous driving (a) Conventional video world models [6, 7, 8, 9, 10] typically serve as data engines for stimulation in pixel space. (b) Unified world models [11, 12, 13] perform video generation and planning as separate tasks. (c) Our proposed policy world model performs planning based on the learned world knowledge.
We evaluate PWM on the popular benchmarks including nuScenes [14], NAVSIM [15]. Notably, our approach achieves strong performance and significantly reduces the average collision rate compared to state-of-the-art methods on nuScenes. Additionally, using only camera inputs, we achieve PDMS performance comparable to state-of-the-art methods that utilize both camera and LiDAR data on the NAVSIM benchmark. These findings both enhance autonomous driving safety and underscore the potential of learning from video-based environmental representations.
@@ 2 Related Work
End-to-end Autonomous Driving. End-to-end autonomous driving has advanced remarkably rapidly. Modern systems [16, 17, 18, 19] map raw sensor inputs directly to control actions, thereby simplifying the traditional perception-to-control pipeline. Early methods [20, 21] used structured bird’s-eye view (BEV) representations, while later work adopted dense 3D occupancy grids [22, 23] or sparse queries [24, 25] for richer, more efficient scene understanding, though handcrafted intermediate representations can limit generalization. More recently, autonomous driving research has begun to incorporate large language models (LLMs) [26, 27] and multimodal LLMs (MLLMs) [28, 29, 30]: driving scenes can be encoded as text for reasoning, achieving strong results in simulators [31, 32, 33], while video-based approaches leverage them to predict actions and answer queries [34, 35, 36, 37].
@@
Generative World Models. World models enable agents to understand and predict future environments from experience [38, 39, 40, 41, 42, 43, 44]. In autonomous driving, various approaches have been proposed for world model construction. Bevworld [45] and WoTE [46] operate in the BEV latent space, while Drive-OccWorld [10] predicts future occupancy states to interact with planning modules. For video generation, generative world models synthesize controllable and high-quality driving scenes, typically through large-scale pretraining. These models fall into two main categories: diffusion-based and autoregressive.Diffusion-based methods include DriveDreamer [4] for controllable video generation, Vista [8] for large-scale conditional synthesis, GLAD [47] for arbitrary-length videos, and DriveDreamer-2 [5] as well as Drive-WM [48] for multi-view generation. Autoregressive models such as GAIA-1/2 [7, 49] unify frames, text, and actions into a single sequence for multi-condition control and long-horizon prediction, while InfinityDrive [50] and DrivingWorld [51] extend to longer and continuous videos. However, token-by-token prediction makes these models computationally expensive and significantly less practical for efficiency-critical downstream tasks.
@@
Unified Generation and Understanding Models. Building on the rapid progress of LLMs and MLLMs, recent research increasingly emphasizes unifying both visual understanding and generative capabilities within a single MLLM [52, 53, 54], often by carefully aligning the representations from large pretrained visual encoders with the embedding space of LLMs. Pioneering unified multimodal models (UMMs) [55, 56, 57] further extend this integration through powerful autoregressive or diffusion-based modeling paradigms. Such holistic unification provides a more principled and scalable framework to externalize the implicit world knowledge encoded in pretrained MLLMs, thereby significantly enhancing their potential for comprehensive world modeling and downstream reasoning.
@@ 3.1 Image Tokenizer with Context-Guided Compression and Decoding
Most driving world models [11, 7, 49] prioritize the quality of the video generation. In order to capture high-resolution details, they typically adopt a long sequence of tokens to represent a single image, leading to significant computational and memory overhead during autoregressive generation. In contrast, PWM focuses on using the learned world representation to benefit planning rather than pursuing photo-realistic details. To achieve a balance between generation quality and efficiency, we employ a new image tokenizer with context-guided compression and decoding.
@@
In our experiment, we adopt a pre-trained image tokenizer [55] as the high-resolution branch $Q_{h}$, which is freezed during training. The low-resolution branch $Q_{l}$ implements a learnable adaptation network mirroring the structure of $Q_{h}$, augmented with additional random initialized cross-attention and downsampling blocks. The encoder of $Q_{l}$ is able to tokenize an input frame of $128\times 224$ resolution into a compact feature map of $4\times 7$, giving rise to $L^{\prime}=28$ tokens per frame. We follow the standard VQ-GAN optimization strategy [58] to train the tokenizer. More implementation details are provided in the supplementary material.
@@ 3.2 Policy World Model
Most existing driving world models either solely focus on world simulation [48, 6] or perform world representation and trajectory planning as two separate tasks with a single model [11, 12, 13]. In contrast, our PWM serves as a more cohesive paradigm, which not only integrates future state forecasting and planning within unified architecture, but also ensures the learned world modeling and forecasting ability can explicitly benefit trajectory planning via a collaborative state-action prediction scheme. This is achieved by the following two unique techniques.
@@
Learning World Modeling from Action-Free Video Generation. Most previous autonomous driving methods learn the video world model via action-conditioned video generation, which is highly dependent on labeled training data. More importantly, they fail to perform any forecasting before the action is predicted. Therefore, when the world model is integrated with the policy model in a unified structure [11, 12, 13], their future forecasting ability cannot be fully explored to benefit planning. To circumvent this issue, we propose to pretrain PWM on action-free video generation (Figure 3 (a)). Formally, given a video clip $\mathbf{I}_{1:N}$, we first transform it into image tokens $\{\mathbf{Z}_{1},\mathbf{D}_{1:N}\}$ using the tokenizer in Sec. 3.1, where $\mathbf{Z}_{1}=\{\mathbf{z}_{1}^{1},\mathbf{z}_{1}^{2},\ldots,\mathbf{z}_{1}^{L}\}$ denotes the contextual tokens of the initial frame, and $\mathbf{D}_{t}=\{\mathbf{d}_{1}^{1},\mathbf{d}_{1}^{2},\ldots,\mathbf{d}_{1}^{L^{\prime}}\}$ denote the compressed tokens of the $t$-th frame. PWM is then trained to generate these video token sequence in an autoregressive manner. Thanks to the token compression scheme, PWM can simultaneously generate all the tokens of a image in-parallel, allowing us to replace the conventional next-token generation with the more efficient next-frame generation scheme (See Figure 3 (b)):
@@ 4.1 Experimental Setups
Datasets. We adopt OpenDV-YouTube dataset [3], nuScenes [14], and NAVSIM [15] as main training datasets. OpenDV-YouTube contains 1747 hours of front-camera video at 10 Hz from 244 cities with scene description generated by BLIP2 [59]. nuScenes includes 1000 scenes (5.5 hours in total) with six camera views per scene, and is annotated with six 3-second waypoints and text descriptions [60]. NAVSIM is built upon OpenScene [61], including 103k training and 12k test samples. We train the tokenizer on the the OpenDV-YouTube. The PWM model is pretrained on OpenDV-YouTube for action-free video generation and then finetuned on nuScenes and NAVSIM for planning. We use only front-view camera video, with 10 Hz for OpenDV-YouTube and NAVSIM, and 12 Hz for nuScenes.
@@
Metrics. To assess the quality and consistency of future video generation, we employ three standard video synthesis metrics: FVD [62], LPIPS [63], and PSNR [64]. For the planning tasks, we use L2 error (m) and collision rate ($\%$) on nuScenes, and Predictive Driver Model Score (PDMS) on NAVSIM. PDMS is a composite metric designed to better align with closed-loop behavior, which combines five sub-scores: no-at-fault collisions (NC), drivable area compliance (DAC), time-to-collision (TTC), comfort (Comf.), and ego progress (EP).
@@
Implementation Details. Our model is initialized from Show-o [55]. During tokenizer training, the frozen branch processes $256\times 448$ images to produce 448 tokens per frame, while the trainable branch uses $128\times 224$ images, generating 28 tokens. Each branch has a separate codebook of size 8192. We sample non-overlapping clips from the OpenDV-YouTube dataset, using 1% for validation and the rest for training. Training runs for $8\times 10^{5}$ steps using AdamW on 6 NVIDIA A800 GPUs with a learning rate of $2\times 10^{-4}$. After training, the tokenizer is frozen. Next, we pre-train the autoregressive world model on OpenDV-YouTube for $3\times 10^{5}$ steps using AdamW with a learning rate of $1\times 10^{-4}$. To prevent catastrophic forgetting, we also conduct auxiliary training on CC12M [65] and FineWeb [66] for image captioning and text generation. For downstream tasks, we fine-tune the model using only the front-view camera. On nuScenes, we condition on 1 second of history to predict 11 frames and 6 waypoints over 3 seconds. Training runs for 16 epochs on 2 A800 GPUs with batch size 8 and learning rate $3\times 10^{-5}$. On NAVSIM, we use 2 seconds of history to predict 10 frames and 8 waypoints across 4 seconds, training for 20 epochs with batch size 14. Dynamic Focal Loss is used throughout. No data augmentation is applied. More details are provided in the Appendix  A.
@@
Table 1: Comparison on the nuScenes validation split. Metrics are computed following the same protocol as [67]. For a fair comparison, results are reported separately for settings without and with ego status (marked with $\dagger$}”); results for UniAD and VAD are reproduced from BEV-Planner [68]. The best results are bolded.
@@
ST-P3 [20] 1.59 2.64 3.73 2.65 0.69 3.62 8.39 4.23
UniAD [21] 0.59 1.01 1.48 1.03 0.16 0.51 1.64 0.77
VAD-Base [67] 0.69 1.22 1.83 1.25 0.06 0.68 2.52 1.09
BEV-Planner [68] 0.30 0.52 0.83 0.55 0.10 0.37 1.30 0.59
Omni-Q [60] 1.15 1.96 2.84 1.98 0.80 3.12 7.46 3.79
LAW [69] 0.24 0.46 0.76 0.49 0.08 0.10 0.39 0.19
Drive-OccWorld [10] 0.25 0.44 0.72 0.47 0.03 0.08 0.22 0.11
@@
VADv2 [18] C&L 97.2 89.1 76.0 91.6 100.0 80.9
TransFuser [71] C&L 97.7 92.8 79.2 92.8 100.0 84.0
DRAMA [72] C&L 98.0 93.1 80.1 94.8 100.0 85.5
Hydra-MDP [19] C&L 98.3 96.0 78.7 94.6 100.0 86.5
DiffusionDrive [17] C&L 98.2 96.2 82.2 94.7 100.0 88.1
UniAD [21] C 97.8 91.9 78.8 92.9 100.0 83.4
LTF [71] C 97.4 92.8 79.0 92.4 100.0 83.8
PARA-Drive [73] C 97.9 92.4 79.3 93.0 99.8 84.0
LAW [69] C 96.4 95.4 81.7 88.7 99.9 84.6
DrivingGPT [11] SC 98.9 90.7 79.7 94.9 95.6 82.4

### v2

@@ 1 Introduction
Driving world models have recently garnered growing research interest, due to their capacity to simulate future environmental states, enabling autonomous systems to anticipate complex traffic dynamics and enhance decision-making safety [survey1, survey2]. This trend is particularly evident in video-based paradigms, driven by the accessibility of large-scale video datasets and the breakthroughs of the video generation technique [genad, drivedreamer, drivedreamer_2]. Despite remarkable progress having been achieved, existing world model based autonomous driving methods  [Adriver-i, Gaia-1, Vista, DrivingDojo, drivinginoccworld] mostly operate in a decoupled manner (Figure 1 (a)), i.e., the world models aim to predict the next state and the associated reward but cannot directly perform trajectory planning. A separate policy model is still required for perceiving and planning. As a result, the full potential of the world model in autonomous driving is largely restricted [Drivinggpt].
Several concurrent works have made the initial attempt towards integrating both world modeling and planning in a unified autoregressive model [doe-1, Drivinggpt], where interleaved image and action token sequences are generated through the next-token prediction manner. Although the two tasks are jointly learned in these works, they are still independently conducted (Figure 1 (b)), where world modeling prioritizes high-fidelity next frame prediction with photo-realistic details conditioned on input actions, and trajectory planning is performed through an end-to-end mapping from the visual observation to output actions without explicitly leveraging the learned world model. As such, the world modeling and trajectory planning are only unified in terms of model architecture and prediction manner, while their synergistic mechanism remains under-explored. It is still unknown whether and how this unification can further benefit autonomous driving.
@@
Figure 1: Comparison of video world models for autonomous driving (a) Conventional video world models [Adriver-i, Gaia-1, Vista, DrivingDojo, drivinginoccworld] typically serve as data engines for stimulation in pixel space. (b) Unified world models [Drivinggpt, doe-1, vavim] perform video generation and planning as separate tasks. (c) Our proposed policy world model performs planning based on the learned world knowledge.
We evaluate PWM on the popular benchmarks including nuScenes [nuScenes], NAVSIM [navsim]. Notably, our approach achieves strong performance and significantly reduces the average collision rate compared to state-of-the-art methods on nuScenes. Additionally, using only camera inputs, we achieve PDMS performance comparable to state-of-the-art methods that utilize both camera and LiDAR data on the NAVSIM benchmark. These findings both enhance autonomous driving safety and underscore the potential of learning from video-based environmental representations.
@@ 2 Related Work
End-to-end Autonomous Driving. End-to-end autonomous driving has advanced remarkably rapidly. Modern systems [ren2024diffusion, diffusiondrive, vadv2, hydra_mdp] map raw sensor inputs directly to control actions, thereby simplifying the traditional perception-to-control pipeline. Early methods [stp3, uniad] used structured bird’s-eye view (BEV) representations, while later work adopted dense 3D occupancy grids [tong2023scene, OccWorld] or sparse queries [sun2024sparsedrive, zhang2024sparsead] for richer, more efficient scene understanding, though handcrafted intermediate representations can limit generalization. More recently, autonomous driving research has begun to incorporate large language models (LLMs) [touvron2023llama, yang2025qwen3] and multimodal LLMs (MLLMs) [liu2023visual, chen2024expanding, bai2025qwen2]: driving scenes can be encoded as text for reasoning, achieving strong results in simulators [LMDrive, zhang2024ad, liu2025dsdrive], while video-based approaches leverage them to predict actions and answer queries [yuan2024rag, DriveLM, DriveGPT4, chen2024driving].
@@
Generative World Models. World models enable agents to understand and predict future environments from experience [ha2018world, dreamerv2, singer2022make, he2022latent, zhou2022magicvideo, khachatryan2023text2video, liu2024sora]. In autonomous driving, various approaches have been proposed for world model construction. Bevworld [bevworld] and WoTE [bevworld2] operate in the BEV latent space, while Drive-OccWorld [drivinginoccworld] predicts future occupancy states to interact with planning modules. For video generation, generative world models synthesize controllable and high-quality driving scenes, typically through large-scale pretraining. These models fall into two main categories: diffusion-based and autoregressive.Diffusion-based methods include DriveDreamer [drivedreamer] for controllable video generation, Vista [Vista] for large-scale conditional synthesis, GLAD [xie2025glad] for arbitrary-length videos, and DriveDreamer-2 [drivedreamer_2] as well as Drive-WM [drive_wm] for multi-view generation. Autoregressive models such as GAIA-1/2 [Gaia-1, Gaia-2] unify frames, text, and actions into a single sequence for multi-condition control and long-horizon prediction, while InfinityDrive [InfinityDrive] and DrivingWorld [hu2024drivingworld] extend to longer and continuous videos. However, token-by-token prediction makes these models computationally expensive and significantly less practical for efficiency-critical downstream tasks.
@@
Unified Generation and Understanding Models. Building on the rapid progress of LLMs and MLLMs, recent research increasingly emphasizes unifying both visual understanding and generative capabilities within a single MLLM [li2025synergen, dong2023dreamllm, tong2024metamorph], often by carefully aligning the representations from large pretrained visual encoders with the embedding space of LLMs. Pioneering unified multimodal models (UMMs) [showo, team2024chameleon, zhou2024transfusion] further extend this integration through powerful autoregressive or diffusion-based modeling paradigms. Such holistic unification provides a more principled and scalable framework to externalize the implicit world knowledge encoded in pretrained MLLMs, thereby significantly enhancing their potential for comprehensive world modeling and downstream reasoning.
@@ 3.1 Image Tokenizer with Context-Guided Compression and Decoding
Most driving world models [Drivinggpt, Gaia-1, Gaia-2] prioritize the quality of the video generation. In order to capture high-resolution details, they typically adopt a long sequence of tokens to represent a single image, leading to significant computational and memory overhead during autoregressive generation. In contrast, PWM focuses on using the learned world representation to benefit planning rather than pursuing photo-realistic details. To achieve a balance between generation quality and efficiency, we employ a new image tokenizer with context-guided compression and decoding.
@@
In our experiment, we adopt a pre-trained image tokenizer [showo] as the high-resolution branch $Q_{h}$, which is freezed during training. The low-resolution branch $Q_{l}$ implements a learnable adaptation network mirroring the structure of $Q_{h}$, augmented with additional random initialized cross-attention and downsampling blocks. The encoder of $Q_{l}$ is able to tokenize an input frame of $128\times 224$ resolution into a compact feature map of $4\times 7$, giving rise to $L^{\prime}=28$ tokens per frame. We follow the standard VQ-GAN optimization strategy [vq-gan] to train the tokenizer. More implementation details are provided in the supplementary material.
@@ 3.2 Policy World Model
Most existing driving world models either solely focus on world simulation [drive_wm, Adriver-i] or perform world representation and trajectory planning as two separate tasks with a single model [Drivinggpt, doe-1, vavim]. In contrast, our PWM serves as a more cohesive paradigm, which not only integrates future state forecasting and planning within unified architecture, but also ensures the learned world modeling and forecasting ability can explicitly benefit trajectory planning via a collaborative state-action prediction scheme. This is achieved by the following two unique techniques.
@@
Learning World Modeling from Action-Free Video Generation. Most previous autonomous driving methods learn the video world model via action-conditioned video generation, which is highly dependent on labeled training data. More importantly, they fail to perform any forecasting before the action is predicted. Therefore, when the world model is integrated with the policy model in a unified structure [Drivinggpt, doe-1, vavim], their future forecasting ability cannot be fully explored to benefit planning. To circumvent this issue, we propose to pretrain PWM on action-free video generation (Figure 3 (a)). Formally, given a video clip $\mathbf{I}_{1:N}$, we first transform it into image tokens $\{\mathbf{Z}_{1},\mathbf{D}_{1:N}\}$ using the tokenizer in Sec. 3.1, where $\mathbf{Z}_{1}=\{\mathbf{z}_{1}^{1},\mathbf{z}_{1}^{2},\ldots,\mathbf{z}_{1}^{L}\}$ denotes the contextual tokens of the initial frame, and $\mathbf{D}_{t}=\{\mathbf{d}_{1}^{1},\mathbf{d}_{1}^{2},\ldots,\mathbf{d}_{1}^{L^{\prime}}\}$ denote the compressed tokens of the $t$-th frame. PWM is then trained to generate these video token sequence in an autoregressive manner. Thanks to the token compression scheme, PWM can simultaneously generate all the tokens of a image in-parallel, allowing us to replace the conventional next-token generation with the more efficient next-frame generation scheme (See Figure 3 (b)):
@@ 4.1 Experimental Setups
Datasets. We adopt OpenDV-YouTube dataset [genad], nuScenes [nuScenes], and NAVSIM [navsim] as main training datasets. OpenDV-YouTube contains 1747 hours of front-camera video at 10 Hz from 244 cities with scene description generated by BLIP2 [blip2]. nuScenes includes 1000 scenes (5.5 hours in total) with six camera views per scene, and is annotated with six 3-second waypoints and text descriptions [ominidrive]. NAVSIM is built upon OpenScene [openscene], including 103k training and 12k test samples. We train the tokenizer on the the OpenDV-YouTube. The PWM model is pretrained on OpenDV-YouTube for action-free video generation and then finetuned on nuScenes and NAVSIM for planning. We use only front-view camera video, with 10 Hz for OpenDV-YouTube and NAVSIM, and 12 Hz for nuScenes.
@@
Metrics. To assess the quality and consistency of future video generation, we employ three standard video synthesis metrics: FVD [fvd], LPIPS [lpips], and PSNR [psnr]. For the planning tasks, we use L2 error (m) and collision rate ($\%$) on nuScenes, and Predictive Driver Model Score (PDMS) on NAVSIM. PDMS is a composite metric designed to better align with closed-loop behavior, which combines five sub-scores: no-at-fault collisions (NC), drivable area compliance (DAC), time-to-collision (TTC), comfort (Comf.), and ego progress (EP).
@@
Implementation Details. Our model is initialized from Show-o [showo]. During tokenizer training, the frozen branch processes $256\times 448$ images to produce 448 tokens per frame, while the trainable branch uses $128\times 224$ images, generating 28 tokens. Each branch has a separate codebook of size 8192. We sample non-overlapping clips from the OpenDV-YouTube dataset, using 1% for validation and the rest for training. Training runs for $8\times 10^{5}$ steps using AdamW on 6 NVIDIA A800 GPUs with a learning rate of $2\times 10^{-4}$. After training, the tokenizer is frozen. Next, we pre-train the autoregressive world model on OpenDV-YouTube for $3\times 10^{5}$ steps using AdamW with a learning rate of $1\times 10^{-4}$. To prevent catastrophic forgetting, we also conduct auxiliary training on CC12M [CC12M] and FineWeb [FineWeb] for image captioning and text generation. For downstream tasks, we fine-tune the model using only the front-view camera. On nuScenes, we condition on 1 second of history to predict 11 frames and 6 waypoints over 3 seconds. Training runs for 16 epochs on 2 A800 GPUs with batch size 8 and learning rate $3\times 10^{-5}$. On NAVSIM, we use 2 seconds of history to predict 10 frames and 8 waypoints across 4 seconds, training for 20 epochs with batch size 14. Dynamic Focal Loss is used throughout. No data augmentation is applied. More details are provided in the Appendix  A.
@@
Table 1: Comparison on the nuScenes validation split. Metrics are computed following the same protocol as [vad]. For a fair comparison, results are reported separately for settings without and with ego status (marked with $\dagger$}”); results for UniAD and VAD are reproduced from BEV-Planner [ego_status]. The best results are bolded.
@@
ST-P3 [stp3] 1.59 2.64 3.73 2.65 0.69 3.62 8.39 4.23
UniAD [uniad] 0.59 1.01 1.48 1.03 0.16 0.51 1.64 0.77
VAD-Base [vad] 0.69 1.22 1.83 1.25 0.06 0.68 2.52 1.09
BEV-Planner [ego_status] 0.30 0.52 0.83 0.55 0.10 0.37 1.30 0.59
Omni-Q [ominidrive] 1.15 1.96 2.84 1.98 0.80 3.12 7.46 3.79
LAW [law] 0.24 0.46 0.76 0.49 0.08 0.10 0.39 0.19
Drive-OccWorld [drivinginoccworld] 0.25 0.44 0.72 0.47 0.03 0.08 0.22 0.11
@@
VADv2 [vadv2] C&L 97.2 89.1 76.0 91.6 100.0 80.9
TransFuser [transfuser] C&L 97.7 92.8 79.2 92.8 100.0 84.0
DRAMA [drama] C&L 98.0 93.1 80.1 94.8 100.0 85.5
Hydra-MDP [hydra_mdp] C&L 98.3 96.0 78.7 94.6 100.0 86.5
DiffusionDrive [diffusiondrive] C&L 98.2 96.2 82.2 94.7 100.0 88.1
UniAD [uniad] C 97.8 91.9 78.8 92.9 100.0 83.4
LTF [transfuser] C 97.4 92.8 79.0 92.4 100.0 83.8
PARA-Drive [para-drive] C 97.9 92.4 79.3 93.0 99.8 84.0
LAW [law] C 96.4 95.4 81.7 88.7 99.9 84.6
DrivingGPT [Drivinggpt] SC 98.9 90.7 79.7 94.9 95.6 82.4